OpenStudy (pokemon23):

Find four consecutive odd integers such that the sum of the first three exceeds the fourth by 18. Find four consecutive odd integers such that the sum of the first three exceeds the fourth by 18. @Mathematics

OpenStudy (anonymous):

9,11,13,15

OpenStudy (anonymous):

x-4 + x-2 + x = x+2 + 18 3x -6 = x + 20 2x = 26 x = 13 therefore 9, 11, 13 & 15

OpenStudy (anonymous):

9, 11, 13, 15